a 78 kg object experiences the following forces:
65 N [w], 72 N [NW], 87 N [E30S]
find the net force acting on it.
1
Expert's answer
2020-03-16T13:04:52-0400
"F=(-65,0)+(-72\cos{45},72\sin{45})+\\+(87\sin{45},-87\cos{45})"

"F=(-54.4,-10.6)N"

"F=\sqrt{54.4^2+10.6^2}=55.4\ N"

"\theta=\arctan{\frac{10.6}{55.4}}=10.8\degree\text{S of W}"